# CodeUltraFeedback: An LLM-as-a-Judge Dataset for Aligning Large Language Models to Coding Preferences

### v1

@@ Abstract
Evaluating the alignment of large language models (LLMs) with user-defined coding preferences is a challenging endeavour that requires assessing intricate textual LLMs’ outputs. By relying on automated metrics and static analysis tools, existing benchmarks fail to assess nuances in user instructions and LLM outputs, highlighting the need for large-scale datasets and benchmarks for LLM preference alignment. In this paper, we introduce CodeUltraFeedback, a preference dataset of 10,000 complex instructions to tune and align LLMs to coding preferences through AI feedback. We generate responses to the instructions using a pool of 14 diverse LLMs, which we then annotate according to their alignment with five coding preferences using the LLM-as-a-Judge approach with GPT-3.5, producing both numerical and textual feedback. We also present CODAL-Bench, a benchmark for assessing LLM alignment with these coding preferences. Our results show that CodeLlama-7B-Instruct, aligned through reinforcement learning from AI feedback (RLAIF) with direct preference optimization (DPO) using CodeUltraFeedback’s AI feedback data, outperforms 34B LLMs on CODAL-Bench, validating the utility of CodeUltraFeedback for preference tuning. Furthermore, we show our DPO-aligned CodeLlama model improves functional correctness on HumanEval+ compared to the unaligned base model. Therefore, our contributions bridge the gap in preference tuning of LLMs for code and set the stage for further advancements in model alignment and RLAIF for code intelligence.
I Introduction
@@
The advent of recent large language models (LLMs) has ushered in a new era of LLMs with high coding capabilities [1, 2, 3, 4, 5, 6], showcasing remarkable performances across a wide range of downstream tasks including code generation [7, 8, 9, 10], code translation [11, 12], bug fixing [13, 6, 14], and more. As the coding abilities of LLMs continue to surge forward, a crucial question emerges: how well do these capabilities align with the expectations of developers, particularly concerning non-functional requirements such as code readability, efficiency, and adherence to best practices?
@@
Current methodologies for fine-tuning and evaluating LLMs widely focus on core capabilities of LLMs, e.g., translating, summarizing, or reviewing code [15, 16, 17, 18, 19, 20, 21, 22], and functional correctness in diverse code generation scenarios [23, 24, 1, 6, 7, 25]. Other works aim at assessing non-functional properties of LLM-generated code, including code quality [26], runtime efficiency [27] and others [28, 29]. However, they generally consist of refining existing benchmarks to evaluate non-functional requirements, often without considering more open coding problems with intricate user instructions. Additionally, the evaluation often relies on automated metrics and external tools based on rigid standards and patterns, overlooking the nuanced complexity of users’ instructions and LLMs’ outputs. Finally, we underline a lack of large-scale datasets tailored for tuning and aligning LLMs to non-functional requirements.
The existing datasets and benchmarks in code intelligence collectively reveal a significant gap in both tuning LLMs and measuring their alignment to non-functional requirements through a more human-centric approach capable of discerning the intricacies of natural and programming languages. Addressing this gap requires a dual approach: developing datasets tailored to tune and align LLMs with non-functional requirements and benchmarks to evaluate their alignment.
@@
In this paper, we introduce two novel contributions: CodeUltraFeedback and CODAL-Bench (“Code Alignment Benchmark”). CodeUltraFeedback is a preference dataset comprising 10,000 complex instructions with 40,000 meticulously curated LLMs responses aligned with five non-functional requirements (or coding preferences [^1] In the rest of this paper, we use the term “coding preferences” instead of “non-functional requirements” to more precisely capture the process of instructing LLMs with explicit preferences, although both terms have similar meanings.): instruction following, code explanation, code complexity and efficiency, code readability, and coding style. The objective of CodeUltraFeedback is to serve as a dataset for preference tuning of LLMs, leveraging recent advancements in LLM alignment, including UltraFeedback [30], reinforcement learning from AI feedback (RLAIF) [31, 32, 33] and LLM-as-a-Judge [34, 35], predicated upon the advanced judging capabilities of LLMs like GPT-3.5 or GPT-4. We adopt an approach auxiliary to UltraFeedback for building our dataset, and start by tagging each instruction with a coding preference. Then, we generate responses to the instructions using four LLMs randomly selected from a pool of 14 LLMs to achieve diversity and consider various writing styles. Finally, we use LLM-as-a-Judge with GPT-3.5 to rate the alignment of the LLMs’ responses with respect to the instruction’s coding preference, resulting in annotations comprising a numerical rating and a textual rationale for the rating. These AI feedbacks can then be leveraged to align LLMs to generate high-quality responses given a coding preference by tuning the LLM through RLAIF using techniques like direct preference optimization (DPO) [36].
@@
Subsequently, we construct CODAL-Bench by selecting a subset of 500 instructions from CodeUltraFeedback. CODAL-Bench serves a different purpose than CodeUltraFeedback and aims to comprehensively assess and compare the alignment of LLMs with coding preferences. We design a rigorous single-answer grading scheme with LLM-as-a-Judge where GPT-3.5-Turbo or GPT-4-Turbo evaluates a single LLM response at a time in a consistent and objective manner. This approach offers a more nuanced evaluation strategy than previous methods that depend on automated metrics and external tools by leveraging the advanced reasoning capabilities of LLMs. Their ability to discern intricacies and nuances in language allows for a more refined and context-sensitive evaluation of how well code aligns with coding preferences in a human-centric fashion.
In the initial part of this paper, we conduct an exploratory analysis of CodeUltraFeedback’s annotations and demonstrate the strong judging capabilities of GPT-3.5-Turbo. GPT-3.5-Turbo effectively recognizes GPT-4-Turbo as a superior model to itself, demonstrating impartiality and fairness in its judgement while also discerning between responses of different quality. Moreover, our initial exploration also reveals a lack of alignment of 12 LLMs, including strong LLMs such as WizardCoder-33B [37] and DeepSeek-Coder-Instruct-33B [3].
In the second part of this paper, we explore preference tuning of a small LLM, CodeLlama-7B-Instruct [2] using CodeUltraFeedback with supervised fine-tuning (SFT) and RLAIF with DPO [36, 38]. The method relies on using the AI feedback data from CodeUltraFeedback, where LLMs’ responses with high and low ratings are selected for tuning and aligning CodeLlama-7B-Instruct to coding preferences with DPO. DPO encourages the LLM to favour highly rated LLMs’ responses, enabling itself to generate more aligned content at inference. Moreover, we implement QLoRA [39] for efficient fine-tuning and show SFT and DPO can be achieved on a single RTX A5000 GPU (24GB).
@@
Our experiments validate the utility of CodeUltraFeedback for preference tuning. We demonstrate that tuning CodeLlama-7B-Instruct with SFT and DPO substantially improves LLM alignment across all coding preferences with high statistical significance on CODAL-Bench, surpassing all larger LLMs, including CodeLlama-34B-Instruct and WizardCoder-33B when using GPT-3.5-Turbo as a judge. Further, we explore LLM alignment on CODAL-Bench using GPT-4-Turbo as a judge, and highlight more mitigated effects of SFT and DPO, albeit CodeLlama-7B-Instruct still outperforms CodeLlama-13B-Instruct and CodeLlama-34B-instruct in this setup. Finally, we show that preference tuning does not hinder the capability of CodeLlama-7B-Instruct in generating functionally correct code. On the contrary, we highlight that CodeLlama-7B-Instruct tuned with SFT and DPO substantially improves Pass@$k$ on HumanEval [1] and HumanEval+ [7]. Therefore, our work establishes strong evidence of the benefits of SFT and DPO to enhance LLM alignment with human-centric coding preferences, demonstrating improvements in both preference alignment and functional correctness. Therefore, our work paves the way for future advancements in LLM training and evaluation, promising a closer alignment with the nuanced expectations of developers.
To summarize, our contributions are the following:
@@
  • We provide CODAL-Bench, a benchmark to evaluate and compare LLM alignment over five coding preferences: instruction following, code explanation, code complexity and efficiency, code readability, and coding style.
  • We validate the utility of CodeUltraFeedback in facilitating LLM alignment of CodeLlama-7B-Instruct, a small LLM, using SFT and DPO.
@@
II CodeUltraFeedback
Figure 1: Overview of CodeUltraFeedback dataset construction procedure. (i) Coding Preferences and Principles: we define five coding preferences: Instruction-following, Code Explanation, Code Complexity and Efficiency, Code Readability, and Coding Style, and 10 principles for each preference to guide LLMs’ generative process. (ii) Magicoder Evol-Instruct Dataset: CodeUltraFeedback is based on a 10k-instruction subset of Magicoder Evol-Instruct dataset [40, 37], where each instruction is tagged with a coding preference. (iii) LLMs Selection: for each instruction, four LLMs are randomly sampled from a diverse pool of 14 LLMs. (iv) Principle-Driven Code Generation: one principle is randomly selected per LLM to guide and align the code generation process with the assigned coding preference. (v) LLM-as-a-Judge Annotation: LLM-as-a-Judge with GPT-3.5-Turbo is used to judge LLMs’ responses according to coding preferences evaluation criteria (see Table I).
@@ II CodeUltraFeedback
II-A Coding Preferences and Principles.
We start by identifying five coding preferences to guide the creation of our dataset and which are essential to evaluate the broader capabilities of LLMs: (1) Instruction-Following is about the strict adherence of the LLM to the instructions provided by users. This preference is foundational for ensuring that LLMs truly follow the user intent and thus provide personalized responses to instructions. (2) Code Explanation emphasizes the generation of clear code with detailed explanations. It underscores the importance of understandable and adaptable code, serving as a bridge between potentially complex programming solutions and the users. (3) Code Complexity and Efficiency preference underlines the LLM capability to generate code optimized for performance in terms of speed and resource utilization. It is another crucial aspect requiring the LLM to carefully balance the speed and resource usage of the solution. (4) Code Readability serves a distinct purpose to the code explanation preference by emphasizing the clarity and understandability of the code itself through its structure, style, and the presence of meaningful documentation and in-line comments. (5) Coding Style preference focuses on the importance of writing code in a manner that not only meets syntactical correctness but also aligns with the idiomatic practices and stylistic norms of the programming language.
We generate 10 principles per preference using ChatGPT following prior work practices [41, 30, 42]. These principles shortly describe the essence of each preference and are leveraged in the Principle-Driven Code Generation step to guide the LLMs generative process. The rationale for establishing 10 principles per preference is to achieve comprehensive and diverse outputs in this code generation step, which we further describe in Section II-D.
@@
II-B Initial Dataset.
To build CodeUltraFeedback, we rely on Magicoder Evol-Instruct dataset [40], an Evol-Instruct [43, 37] version of CodeAlpaca dataset comprising complex coding problem instructions. We select a random subset of Magicoder Evol-Instruct comprising 10,000 samples serving as an initial dataset for CodeUltraFeedback. We assign 2,000 samples per coding preference, ensuring a balanced representation of each preference in the dataset. The reason for subsampling the dataset to 10,000 samples is to lower costs related to OpenAI API needed to generate some of the responses and the annotations.
TABLE I: Code readability assessment template.
Evaluate the readability of code segments. Assess how comments and documentation contribute to understanding the code’s logic, purpose, and operation.
Evaluation Criteria:
@@
II-C LLMs Selection.
In order to generate highly diverse solutions to the instructions, we select a diverse pool of 14 LLMs spanning eight model families. These include GPT-4-Turbo and GPT-3.5-Turbo as closed-source LLMs. We include Llama-2-13B-Chat and Llama-2-70B-Chat [5], specialized for chat use cases. We choose CodeLlama-7/13/34B-Instruct models [2], instruction-tuned for code generation. We also include WizardLM-15/33B models [43] and Mistral-7B-Instruct [4], which are instruction-tuned LLMs that can follow complex instructions. Finally, we include WizardCoder-15/33B [37] and DeepSeek-Coder-6.7/33B-Instruct [3], which are instruction-tuned LLMs that demonstrate state-of-the-art performance on code generation benchmarks such as HumanEval [1] and HumanEval+ [7]. From this pool, we randomly select four LLMs for each instruction to generate responses, allowing for the representation of a variety of coding approaches and styles.
TABLE II: Excerpt of ratings and rationales generated by GPT-3.5-Turbo using LLM-as-a-Judge prompting. GPT-3.5-Turbo generates clear rationales for each rating demonstrating its strong judging capabilities.
@@
II-D Principle-Driven Code Generation.
In this step, we leverage the concept of principle-driven generation [41, 44, 42] for code. One principle, corresponding to the coding preference of the sample, is randomly chosen per LLM and appended to the instruction in an input prompt to guide the generation. This strategy helps the LLMs generate responses that potentially align with the requirements of the coding preferences. The output of this step is the four LLMs’ responses to the principle-augmented instructions.
II-E LLM-as-a-Judge Annotation.
We apply the LLM-as-a-Judge methodology [34] to annotate responses using GPT-3.5-Turbo. Drawing from practices established in previous work [30], we implement a set of guidelines ensuring GPT-3.5-Turbo delivers both consistent and objective feedback. Each annotation consists of (1) a score on a 1–5 scale offering a comprehensive assessment of the response’s alignment, and (2) a rationale for the assigned score, highlighting the response’s strengths and identifying any potential shortcomings. Tasking GPT-3.5-Turbo to provide a rationale for the scores encourages a chain-of-thought [45] process during the annotation process, elucidating the reasons behind the scores and meaningful feedback. In addition, this process mitigates the risk of variability and subjectivity that might arise from purely quantitative assessments. Consequently, this reflective process guarantees that the scoring is not just a surface-level judgment but can be backed by thoughtful consideration of some well-defined coding preferences evaluation criteria.
Table I illustrates the assessment template and evaluation criteria for code readability evaluation. It aims to provide precise and comprehensive criteria to the LLM judge for response assessment. We design analogous assessment templates, each tailored with specific evaluation criteria, for evaluating the other coding preferences. The annotation procedure entails constructing a prompt combining the coding preference assessment template and the four LLMs responses. GPT-3.5-Turbo is then tasked to generate the annotations for all four responses simultaneously, enhancing the consistency and reliability of the evaluation process over individual response assessments. Table II presents an example of the output of the annotation process using GPT-3.5-Turbo, with the LLMs responses and some rationales omitted for brevity. This example showcases GPT-3.5-Turbo’s capability to provide comprehensive justifications for each rating, demonstrating its effective evaluation capabilities. Furthermore, Table II illustrates an example of relatively intricate instructions included in CodeUltraFeedback, showcasing the complexity of the task. More details about CodeUltraFeedback and the prompt templates are included in our replication package: https://github.com/martin-wey/CodeUltraFeedback.
III Evaluating LLMs on CodeUltraFeedback
TABLE III: Average scores across coding preferences for each LLM. The statistical significance of differences in ratings between each LLM and GPT-3.5-Turbo was tested using pairwise Welch’s $t$-tests with 100,000 permutations to control for type I error. Significance levels are denoted as follows: * for $p<0.05$, $\dagger$ for $p<0.01$, and $\ddagger$ for $p<0.001$. The results from almost all pairwise $t$-tests indicate highly significant differences ($p<0.001$).

### v2

@@ Abstract.
Evaluating the alignment of large language models (LLMs) with user-defined coding preferences is a challenging endeavour that requires a deep assessment of LLMs’ outputs. Existing methods and benchmarks rely primarily on automated metrics and static analysis tools, which often fail to capture the nuances of user instructions and LLM outputs. To address this gap, we propose using the LLM-as-a-Judge methodology to evaluate the alignment of LLMs with coding preferences. Based on this approach, we present CodeUltraFeedback, a comprehensive dataset designed to facilitate the evaluation and improvement of LLM alignment. CodeUltraFeedback consists of 10,000 coding instructions, each annotated with four responses generated from a diverse pool of 14 LLMs. These responses are ranked based on five distinct coding preferences using GPT-3.5 as a judge, providing both numerical scores and detailed textual feedback. Our analysis of CodeUltraFeedback reveals that responses from GPT-3.5 and GPT-4 are generally preferred over those from open-weight LLMs, highlighting significant differences in alignment between closed and open-weight models. In turn, we explore the usage of CodeUltraFeedback as feedback data to fine-tune and align CodeLlama-7B-Instruct using supervised fine-tuning (SFT) and reinforcement learning from AI feedback (RLAIF) with direct preference optimization (DPO). The resulting aligned CodeLlama-7B-Instruct model outperforms larger LLMs in terms of alignment with coding preferences and shows improved functional correctness on the HumanEval+ benchmark compared to the original instruct model. Therefore, our contributions bridge the gap in preference tuning of LLMs for code and set the stage for further advancements in model alignment and RLAIF in automated software engineering.
Large language models, code generation, automated software engineering, reinforcement learning from AI feedback, direct preference optimization
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
The advent of recent large language models (LLMs) has ushered in a new era of LLMs with high coding capabilities (Chen and others., 2021; Roziere et al., 2023; Guo et al., 2024; Jiang et al., 2023; Touvron et al., 2023; Muennighoff et al., 2023), showcasing remarkable performances across a wide range of downstream tasks including code generation (Liu et al., 2023b; Athiwaratkun et al., 2022; Weyssow et al., 2023; Zheng et al., 2023b), code translation (Jiao et al., 2023; Pan et al., 2023), bug fixing (Silva et al., 2023; Muennighoff et al., 2023; Ye et al., 2021), and more. However, as LLMs continue to advance, a critical question emerges: How well do these capabilities align with developers’ expectations, particularly regarding non-functional requirements such as code readability, efficiency, and adherence to best practices?
@@
Current methodologies for fine-tuning and evaluating LLMs primarily focus on core capabilities, e.g., translating, summarizing, or reviewing code (Lu et al., 2021; Puri et al., 2021; Zhu et al., 2022; Niu et al., 2023; Zhou et al., 2023; Sghaier and Sahraoui, 2024; Sghaier et al., 2023; Sghaier and Sahraoui, 2023), and functional correctness in diverse code generation scenarios (Hendrycks et al., 2021; Austin et al., 2021; Chen and others., 2021; Muennighoff et al., 2023; Liu et al., 2023b; Khan et al., 2023). While some research has started to address non-functional properties of LLM-generated code, such as code quality (Siddiq et al., 2023), runtime efficiency (Huang et al., 2024), and other aspects (Yetiştiren et al., 2023; Singhal et al., 2024), these efforts often fall short in several key areas. First, existing benchmarks typically refine established metrics without considering more open-ended coding problems that involve complex user instructions. Second, the assessment of LLM outputs often relies on automated metrics and external tools based on inflexible standards and patterns. This approach fails to capture the nuanced complexity of users’ instructions and the subtleties in LLMs’ responses. Third, there is a notable absence of large-scale datasets specifically designed for tuning and aligning LLMs to non-functional requirements.
The landscape of existing datasets and benchmarks in automated software engineering reveals a critical gap: the lack of methods for both tuning LLMs and accurately measuring their alignment with non-functional requirements. Furthermore, current evaluation methods fall short of capturing the nuances in natural and programming languages, underscoring the need for a more sophisticated, language-centric approach to assessment. To address these shortcomings, we must develop comprehensive datasets designed to improve LLM alignment with non-functional requirements and design evaluation methods capable of evaluating LLM alignment with greater precision. By closing these gaps, we can better bridge the divide between developers’ expectations and the output of LLMs, ensuring that the generated code not only functions correctly but also adheres to best practices, maintains readability, and operates efficiently.
@@
In this paper, we introduce two novel contributions: CodeUltraFeedback and CodeUltraFeedback-Bench. CodeUltraFeedback is a preference dataset comprising 10,000 complex instructions with 40,000 meticulously curated LLMs responses aligned with five non-functional requirements (or coding preferences [^1] In the rest of this paper, we use the term “coding preferences” instead of “non-functional requirements” to more precisely capture the process of instructing LLMs with explicit preferences, although both terms bear similar meanings.): instruction following, code explanation, code complexity and efficiency, code readability, and coding style. The objective of CodeUltraFeedback is to serve as a dataset for preference tuning of LLMs, leveraging recent advancements in LLM alignment, including UltraFeedback (Cui et al., 2023), reinforcement learning from AI feedback (RLAIF) (Bai et al., 2022; Lee et al., 2023; Ouyang et al., 2022) and LLM-as-a-Judge (Zheng et al., 2023a; Li et al., 2023a), predicated upon the advanced judging capabilities of LLMs like GPT-3.5 or GPT-4. We adopt an approach auxiliary to UltraFeedback for building our dataset, starting by tagging each instruction with a coding preference. Then, we generate responses to the instructions using four LLMs randomly selected from a pool of 14 LLMs to achieve diversity and consider various writing styles. Finally, we use LLM-as-a-Judge with GPT-3.5 to rate the alignment of the LLM responses with the instruction’s coding preference, resulting in annotations comprising a numerical rating and a textual rationale for the rating. The resulting dataset can then be leveraged to align an LLM to generate high-quality responses given a coding preference by tuning the LLM through RLAIF using techniques like direct preference optimization (DPO) (Rafailov et al., 2023).
@@
Subsequently, we construct CodeUltraFeedback-Bench by selecting a subset of 500 instructions from CodeUltraFeedback. CodeUltraFeedback-Bench serves a different purpose than CodeUltraFeedback and aims to comprehensively assess and compare the alignment of LLMs with coding preferences. We design a rigorous single-answer grading scheme using LLM-as-a-Judge that allows the judge LLM to evaluate LLM-generated code consistently and objectively. This approach offers a more nuanced evaluation strategy than previous methods that rely on automated metrics and external tools by leveraging the advanced reasoning capabilities of LLMs. Their ability to discern nuances in language allows for a more refined and context-sensitive evaluation of how well LLM-generated code aligns with coding preferences in a language-centric fashion.
@@
An initial data analysis of CodeUltraFeedback’s annotations reveals a lack of alignment of 12 LLMs, including strong LLMs such as WizardCoder-33B (Luo et al., 2023) and DeepSeek-Coder-Instruct-33B (Guo et al., 2024), while GPT-3.5 and GPT-4 demonstrate strong alignment. In turn, we explore preference tuning of a small model, CodeLlama-7B-Instruct (Roziere et al., 2023) using CodeUltraFeedback with supervised fine-tuning (SFT) and RLAIF with DPO (Rafailov et al., 2023; Tunstall et al., 2023). This method selects high and low-rated responses to align CodeLlama-7B-Instruct with coding preferences using DPO, encouraging the model to favor highly rated responses. We demonstrate that tuning CodeLlama-7B-Instruct with SFT and DPO substantially improves alignment across all coding preferences, surpassing all larger LLMs, including CodeLlama-34B-Instruct and WizardCoder-33B when using GPT-3.5-Turbo as a judge. Further, we explore LLM alignment on CodeUltraFeedback-Bench using GPT-4-Turbo as a judge and highlight more mitigated effects of SFT and DPO, albeit CodeLlama-7B-Instruct still outperforms CodeLlama-13B-Instruct and CodeLlama-34B-instruct in this setup. We also show that preference tuning does not hinder CodeLlama-7B-Instruct’s ability to generate functionally correct code, substantially improving Pass@$k$ on HumanEval (Chen and others., 2021) and HumanEval+ (Liu et al., 2023b). Finally, we implement QLoRA (Dettmers et al., 2023) for efficient fine-tuning, showing that SFT and DPO can effectively operate on a single RTX A5000 GPU (24GB). Our work establishes strong evidence of the benefits of SFT and DPO in improving LLM alignment with coding preferences, demonstrating advancements in both preference alignment and functional correctness. This paves the way for future improvements in LLM training and evaluation, promising closer alignment with developers’ nuanced expectations.
To summarize, our main contributions are the following:
@@
  • We contribute CodeUltraFeedback-Bench, a benchmark to evaluate and compare LLM alignment using LLM-as-a-Judge over five coding preferences: instruction following, code explanation, code complexity and efficiency, code readability, and coding style.
  • We validate the utility of CodeUltraFeedback in facilitating LLM alignment of CodeLlama-7B-Instruct using SFT and DPO.
@@
2. CodeUltraFeedback
@@
2.1. Coding Preferences and Principles.
We start by identifying five coding preferences to guide the creation of our dataset and which are essential to evaluate the broader capabilities of LLMs: (1) Instruction-Following is about the strict adherence of the LLM to the instructions provided by users. This preference is foundational for ensuring that LLMs truly follow the user intent and thus provide personalized responses to instructions. (2) Code Explanation emphasizes generating code with detailed natural language explanations. It underscores the importance of an LLM in providing a solution with explanations that can serve as a bridge between potentially complex code and users while improving trustworthiness. (3) Code Complexity and Efficiency preference underlines the LLM capability to generate code optimized for performance in terms of speed and resource utilization. It is another crucial aspect requiring the LLM to balance the speed and resource usage of the solution carefully. (4) Code Readability serves a distinct purpose to the code explanation preference by emphasizing the clarity and understandability of the code itself through its structure, style, and the presence of meaningful documentation and in-line comments. (5) Coding Style preference focuses on the importance of writing code that not only meets syntactical correctness but also aligns with the idiomatic practices and stylistic norms of the programming language.
@@
Figure 1. Overview of CodeUltraFeedback dataset construction procedure. (i) Coding Preferences and Principles: we define five coding preferences: Instruction-following, Code Explanation, Code Complexity and Efficiency, Code Readability, and Coding Style, and 10 principles for each preference to guide LLMs’ generative process. (ii) Magicoder Evol-Instruct Dataset: CodeUltraFeedback is based on a 10k-instruction subset of Magicoder Evol-Instruct dataset (Wei et al., 2023; Luo et al., 2023), where each instruction is tagged with a coding preference. (iii) LLMs Selection: for each instruction, four LLMs are randomly sampled from a diverse pool of 14 LLMs. (iv) Principle-Driven Code Generation: one principle is randomly selected per LLM to guide and align the code generation process with the assigned coding preference. (v) LLM-as-a-Judge Annotation: LLM-as-a-Judge with GPT-3.5-Turbo is used to judge LLMs’ responses according to coding preferences evaluation criteria (see Table 1).
2.2. Initial Dataset.
To build CodeUltraFeedback, we rely on Magicoder Evol-Instruct dataset (Wei et al., 2023), an Evol-Instruct (Xu et al., 2023; Luo et al., 2023) version of CodeAlpaca dataset comprising coding problem instructions. We select a random subset of Magicoder Evol-Instruct comprising 10,000 samples serving as an initial dataset for CodeUltraFeedback. We assign 2,000 samples per coding preference, ensuring a balanced representation of each preference in the dataset. The reason for subsampling the dataset to 10,000 samples is to lower costs related to OpenAI API needed to generate some of the responses and the annotations.
2.3. LLMs Selection.
In order to generate highly diverse solutions to the instructions, we select a pool of 14 LLMs spanning eight model families. These include GPT-4-Turbo and GPT-3.5-Turbo as closed-source LLMs. We include Llama-2-13B-Chat and Llama-2-70B-Chat (Touvron et al., 2023), specialized for chat use cases. We choose CodeLlama-7/13/34B-Instruct models (Roziere et al., 2023), instruction-tuned for code generation. We also include WizardLM-15/33B models (Xu et al., 2023) and Mistral-7B-Instruct (Jiang et al., 2023), which are instruction-tuned LLMs that can follow complex instructions. Finally, we include WizardCoder-15/33B (Luo et al., 2023) and DeepSeek-Coder-6.7/33B-Instruct (Guo et al., 2024), recent instruction-tuned LLMs that demonstrate state-of-the-art performance on code generation benchmarks such as HumanEval (Chen and others., 2021) and HumanEval+ (Liu et al., 2023b). From this pool, we randomly select four LLMs for each instruction to generate responses, allowing for the representation of a variety of coding approaches and styles.
2.4. Principle-Driven Code Generation.
In this step, we leverage principle-driven generation and generate 10 principles per coding preference using ChatGPT following prior work practices (Sun et al., 2023; Cui et al., 2023; Mukherjee et al., 2023). The rationale for establishing 10 principles per preference is to diversify the dataset. All principles for each coding preference can be found in Tables 1620.
During code generation, one principle, corresponding to the coding preference of the sample, is randomly chosen per LLM and appended to the instruction in an input prompt to guide the generation. This strategy helps the LLMs generate responses that potentially align with the requirements of the coding preferences. The output of this step is the four LLMs’ responses to the principle-augmented instructions.
2.5. LLM-as-a-Judge Annotation.
Table 1. Code readability assessment template.
Evaluate the readability of code segments. Assess how comments and documentation contribute to understanding the code’s logic, purpose, and operation. Evaluation Criteria:
@@
We apply the LLM-as-a-Judge methodology (Zheng et al., 2023a) to annotate responses using GPT-3.5-Turbo. Drawing from practices established in previous work (Cui et al., 2023), we implement a set of guidelines ensuring GPT-3.5-Turbo delivers both consistent and objective feedback. Each annotation consists of (1) a score on a 1–5 scale offering a comprehensive assessment of the response’s alignment, and (2) a rationale for the assigned score, highlighting the response’s strengths and identifying any potential shortcomings. Tasking GPT-3.5-Turbo to provide a rationale for the scores encourages a chain-of-thought (Wei et al., 2022) process during the generation of annotations, elucidating the reasons behind the scores and meaningful feedback. In addition, this process mitigates the risk of variability and subjectivity that might arise from purely quantitative assessments. Consequently, this reflective process guarantees that the scoring is not just a surface-level judgment but can be backed by thoughtful consideration of some well-defined coding preferences evaluation criteria.
Table 1 illustrates the assessment template and evaluation criteria for code readability evaluation. It aims to provide precise and comprehensive criteria to the LLM judge for response assessment. We design analogous assessment templates, each tailored with specific evaluation criteria, for evaluating the other coding preferences (see Tables 69). The annotation procedure entails constructing a prompt combining the coding preference assessment template and the four LLMs responses. GPT-3.5-Turbo is then tasked to generate the annotations for all four responses simultaneously, enhancing the consistency and reliability of the evaluation process over individual response assessments. Table 2 presents an example of the output of the annotation process using GPT-3.5-Turbo, with the LLMs responses and some rationales omitted for brevity. This example showcases GPT-3.5-Turbo’s capability to provide comprehensive justifications for each rating, demonstrating its effective evaluation capabilities. Furthermore, Table 2 illustrates an example of relatively complex instructions included in CodeUltraFeedback, showcasing the complexity of the task.
Table 2. Excerpt of ratings and rationales generated by GPT-3.5-Turbo using LLM-as-a-Judge prompting. GPT-3.5-Turbo generates clear rationales for each rating demonstrating its strong judging capabilities.
@@
3. CodeUltraFeedback: Dataset Analysis
Table 3. Average scores across coding preferences for each LLM. The statistical significance of differences in ratings between each LLM and GPT-3.5-Turbo was tested using pairwise Welch’s $t$-tests. Significance levels are denoted as follows: * for $p<0.05$, $\dagger$ for $p<0.01$, and $\ddagger$ for $p<0.001$. The results from almost all pairwise $t$-tests indicate highly significant differences ($p<0.001$).